OpenStudy (hatahoooochee):

Help, I'll fan and medal Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = Four divided by x squared. + 9

OpenStudy (hatahoooochee):

Answers: f(x) = x + 9, g(x) = Four divided by x squared. f(x) = x, g(x) = Four divided by x. + 9 f(x) = One divided by x., g(x) = Four divided by x. + 9 f(x) = Four divided by x squared., g(x) = 9

OpenStudy (chrishogan):

Use this, y=(4/x^2)+9

OpenStudy (hatahoooochee):

I'm confused

OpenStudy (chrishogan):

How so?

OpenStudy (hatahoooochee):

I just don't understand what I'm supposed to do

OpenStudy (hatahoooochee):

@Vocaloid

OpenStudy (hatahoooochee):

@mayankdevnani